OpenStudy (shamil98):

Implicit Differentiation, just need my work checked.. x^2y + xy^2 = 6

OpenStudy (raffle_snaffle):

Are you taking respect to x or y?

OpenStudy (shamil98):

Respect to y. \[\large x^2y + xy^2 = 6\]

OpenStudy (raffle_snaffle):

so 

OpenStudy (shamil98):

\[\large \frac{ d }{ dx } x^2y + \frac{ d }{ dx } xy^2 = \frac{ d }{ dx } 6\] Product rule. 2xy + x^2 + y^2\[\large 2xy + x^2y' + y^2+ x2yy' = 0\] Is this correct so far?...

OpenStudy (raffle_snaffle):

oops i meant